forty eight teachers from a certain region were surveyed. The contingency table​ cross-classifies these teachers by highest degree obtained and whether they offer field trips.

Field Trips

Yes left parenthesis Upper F 1 right parenthesisF1

No left parenthesis Upper F 2 right parenthesisF2

Total

​Bachelor's left parenthesis Upper D 1 right parenthesisD1

1313

1313

2626

​Master's left parenthesis Upper D 2 right parenthesisD2

1616

66

2222

Total

2929

1919

4848

a. How many of these teachers offered field​ trips?

There are

nothing

teachers who offered field trips.

b. How many of these teachers have​ master's degrees?

There are

nothing

teachers with a​ master's degree.

c. How many teachers with only​ bachelor's degrees offered field​ trips?

There are

nothing

teachers with only a​ bachelor's degree who offered field trips.d. Compute the probability of event

Upper D 1D1.

Upper P font size decreased by 5 left parenthesis Upper D 1 right parenthesisP D1equals=nothing

​(Round to three decimal places as​ needed.)Compute the probability of event

left parenthesis Upper D 2 & Upper F 2 right parenthesisD2 & F2.

Upper P font size decreased by 5 left parenthesis Upper D 2 & Upper F 2 right parenthesisP D2 & F2equals=nothing

​(R
Answer #1

a. How many of these teachers offered field​ trips

29

b. How many of these teachers have​ master's degrees

22

c. How many teachers with only​ bachelor's degrees offered field​ trips

13

d. Compute the probability of event Upper D 1

P(D1) = 26/48 = 0.5417

e .

P(D2&F2) = 6/48 = 0.125